OpenStudy (anonymous):

a sphere has a volume of 322cm3 what is the sphere's surface area? Could someone walk me through step by step please

OpenStudy (callisto):

Volume of a sphere \[=\frac{4}{3}\pi r^3\] It is given that the volume is 322cm^3, so you can express the radius r in term of \(\pi\) Then, find the surface area by using the formula Surface area of a sphere \[=4\pi r^2\]

OpenStudy (anonymous):

how do you express it in Pi and how do i find the radius?

OpenStudy (callisto):

\[322= \frac{4}{3}\pi r^3\] Divide both sides by 4/3, what do you get?

OpenStudy (anonymous):

241.5Pi^3 ???

OpenStudy (callisto):

Hmm.. Not really.. \[241.5 = \pi r^3\]\[r^3=\frac{241.5}{\pi}\]\[r=(\frac{241.5}{\pi})^\frac{1}{3}\] Got it?

OpenStudy (anonymous):

is the radius 25.6?

OpenStudy (callisto):

No...

OpenStudy (anonymous):

4.25?

OpenStudy (callisto):

Yes :) Now, you can calculate the surface area :)

OpenStudy (anonymous):

we got 25.6 how did you get 4.25

OpenStudy (callisto):

\(x^\frac{1}{3} \) is \(\sqrt[3]{x}\) , not \(\frac{x}{3}\)

OpenStudy (anonymous):

ok thanks i think i got it :)

OpenStudy (callisto):

You're welcome :)